OpenStudy (flvs.net):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 3) is on the graphed line. Part A: What is the price of 1 pound of cabbage? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (10, 6) on the graph represent? Explain in words. (2 points)

OpenStudy (flvs.net):

@Michele_Laino

OpenStudy (flvs.net):

Can you please help me??

OpenStudy (michele_laino):

I need to know the graph mentioned into your exercise

OpenStudy (flvs.net):

There isn't one. That's why I need help.

OpenStudy (michele_laino):

we have 3 dollars for 5 pounds of cabbage, so the unit price, is: 3/5=...

OpenStudy (flvs.net):

Okay.

OpenStudy (flvs.net):

Is that it??

OpenStudy (michele_laino):

yes! the requested proportion is: x:1=3:5

OpenStudy (flvs.net):

Okay, thank you!

OpenStudy (michele_laino):

we have to do part B

OpenStudy (flvs.net):

Yes

OpenStudy (michele_laino):

the ordere pair (6,10) represents the price of 6 dollars for 10 pounds of cabbage

OpenStudy (michele_laino):

ordered*

OpenStudy (flvs.net):

Wait so what's the answer for A? @Michele_Laino

OpenStudy (michele_laino):

the answer fo part A is given by the subsequent unit price: x= 3/5=... with this proportion: x:1=3:5

OpenStudy (michele_laino):

what is 3/5 rounded to the nearest hundredth?

OpenStudy (flvs.net):

idk

OpenStudy (flvs.net):

:(

OpenStudy (flvs.net):

Are you there @Michele_Laino

OpenStudy (michele_laino):

it is simple: 3/5=0.60

OpenStudy (flvs.net):

Thank you so much!

OpenStudy (michele_laino):

:)